# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@ 6 Data Perspective
In this section, we analyze current work from a data perspective, including time series and spatio-temporal data. We focus on introducing how diffusion models capture the unique properties of different data modalities.
6.1 Time Series Data
@@
6.1.1 Univariate Time Series
@@
Univariate time series is characterized by having only one variable of interest observed over a period of time. Data in this category include ECG signal, audio, electricity load, etc. The complex sequential patterns (e.g., the trend and the periodicity) are the core data property in univariate time series, which pose great challenges in developing relevant diffusion models.
@@
Diffusion Model for Univariate Time Series. Diffusion models for univariate time series are primarily developed to model the uncertainty present in the data, facilitating tasks such as probabilistic forecasting or data generation. An essential element within the diffusion model is the denoising network, which determines the extent of noise removal at each stage. In the context of univariate time series, the denoising network commonly employs the 1-D CNN to identify sequential patterns in the input data. For instance, both WaveGrad [20] and DiffWave [91] incorporate a 1-D CNN in the denoising network to extract local sequential features from audio sequences. Furthermore, some studies opt to transform the univariate time series into the frequency domain initially to enhance the capture of long-term sequential correlations from a global standpoint. This transformation converts the time series from 1-D to 2-D data, enabling the use of a 2-D CNN in the denoising network to capture correlations within the frequency domain, such as [47] for the stock price forecasting and DiffLoad [28] for electricity load forecasting.
6.1.2 Multivariate Time Series
@@
Multivariate time series is characterized by having multiple variables of interest observed over the same period of time. Instead of the sequential patterns in each variate, multivariate time series also leverage the interdependencies among different variables to capture more comprehensive information downstream tasks.
@@
Diffusion Models for Multivaritate Time Series. Beyond approaches that focus on univariate time series, there have been efforts towards multivariate time series. Multivariate time series is naturally a 2-D data similar to the image data, where the signal of each variate can be considered as one channel in the image. To this end, diffusion models for multivariate time series usually adopt the vanilla Unet as the denoising net, which is a common practice in diffusion models for images.
@@
For example, [5] first transforms demultiplex, denoising, and interpolation into image-to-image transformation tasks, then leverages the diffusion model for different tasks. In the demultiplex task, the diffusion model learns to remove the multiples without removing primary energy. In the denoising and interpolation task, the diffusion model learns to eliminate undesired uncorrelated noise and recover the missing values, while preserving the inherent characteristics of the data. MIDM [262] proposes a novel multivariate imputation diffusion model that incorporates correlations between observed and missing values. By re-deriving the ELBO of the conditional diffusion model, MIDM ensures consistency between observed and missing data, thus leading to improved imputation accuracy. Diff-E [109] further utilizes the diffusion model for representation learning, which combines DDPMs with a conditional autoencoder to enhance the decoding performance of speech-related EEG signals.
6.2 Spatio-Temporal Data
@@
In real-world systems, a myriad of elements interact with each other both spatially and temporally, resulting in a spatio-temporal composition. Spatio-temporal data (STD) is the de facto most popular data structure for injecting such structural information into the formulation of practical problems. In this section, we introduce developments of diffusion models for spatial-temporal data, mainly with two data modalities: (1) the Spatio-Temporal Graph (STG) and (2) the Spatio-Temporal Trajectory (STT). Unlike diffusion models for time series data, existing works in STD need to model the temporal dependencies as well as the spatial dependencies, making the task even more challenging.
6.2.1 Spatio-Temporal Graph
@@
The spatio-temporal graph is generated by sensors at different places in a period of time, where the correlation of those sensors is typically described as a graph. We categorize diffusion models for spatial-temporal graph into domain-oriented and domain-agnostic works.
@@
Most of the existing works fall into the category of domain-oriented spatial-temporal graph diffusion models. They typically leverage the powerful distribution learning abilities of diffusion models for spatial-temporal data mining tasks in specific domains, such as traffic and climate. In this research trend, the majority of works come from the traffic domain. DiffUFlow [35] represents pioneering efforts as traffic diffusion models, which convert the fine-grained urban flow inference as a denoising diffusion process. SpecSTG [88] advances the field by conducting the diffusion process in the spectral space projected by the STG, resulting in faster inference speeds. In a similar vein, [59] uses a diffusion model to infer the link probability and reconstruct causal graphs in the decoder stage adaptively for the STG forecasting task. Furthermore, [65] solves the origin-destination travel time estimation problem with a two-step process, where the first step predicts the possible travel route with a diffusion model conditioned on the given origin-destination pair. Beyond the traffic domain, there have been initiatives towards the climate. A notable example is [78], which quantifies the uncertainty of air quality forecasting based on the spatial-temporal diffusion model. Similarly, SRNDiff [85] conducts the short-term rainfall nowcasting through the conditional diffusion model.
@@
In the second group, researchers focus on developing domain-agnostic models, which can achieve promising performance across a variety of domains. For instance, DiffSTG represents a pioneering effort by introducing a unified diffusion framework for multiple STG tasks such as forecasting and imputation. Concurrently, PriSTI [111] and [69] study the spatial-temporal imputation task while treating the task of recovering unobserved values as a denoising process conditioned on the observed values. Similarly, DYffusion [29] models the temporal dynamics directly within diffusion steps, leading to a stochastic, time conditioned forecasting network. Another noteworthy contribution is [170] under the domain-agnostic and task-agnostic setting, which aims to harness the power of diffusion models, and unify diffusion models for probabilistic spatio-temproal graph learning.
6.2.2 Spatio-Temporal Trajectory
@@
Spatial-temporal trajectory is a sequence of locations ordered by time that describe the movements of an object in a geographical space. Analysis of spatial-temporal modality is particularly crucial to discovering the mobility patterns of moving objects, which serves as the foundation for many downstream tasks, such as POI recommendation and next-location forecasting. In terms of diffusion models for trajectory data, most works are developed to solve the trajectory generation task. They leverage the ability of the diffusion model to learn high-dimensional data distribution while injecting spatial-temporal correlation into the diffusion process. As exemplified by DiffTraj [81], which generates the GPS trajectory with the diffusion probabilistic model in an unconditioned manner. Similarly, [60] utilizes the diffusion model for generating high-quality human mobility data. More recently, Diff-RNTraj [17] further generates the trajectory constrained by the road-network. Beyond generation, there is a growing interest in applying diffusion models for trajectory forecasting tasks, and [66] and [184] exemplify this trend.
7 Application Perspective
@@
In this section, we summarize and discuss the most important applications of diffusion models in time series and spatio-temporal data, including healthcare, smart city, recommendation, climate and weather, energy and electricity, audio, video and so on. The toy signal examples of each application are shown in  Figure 8 .
@@
Figure 8: Examples of time series and spatio-temporal data for different application scenarios.
TABLE II: Summary of dataset resources in different application domains.
Applications Dataset Description Timeframe Source Citations
Healthcare Seed V dataset From 16 participants in three emotions test sessions through video presentation. Each test is around 50 minutes. [118] [119]
QT Database 105 fifteen-minute excerpts of two-channel ECG Holter recordings. Examples of each morphology were included in this subset of annotated beats; at least 30 beats in each record, 3622 beats in all, were manually annotated in the database. 105*15 minutes [120] [57]
MIT-BIH NST dataset 12 half-hour ECG recordings and 3 half-hour recordings of noise typical in ambulatory ECG recordings. 15 half hour [121] [57]
PTB-XL dataset Collection of clinical 12-lead ECG data comprising 21,837 records from 18,885 patients. Each 10 seconds length. [122] [123, 124]
MIT-BIH dataset Healthy with 90,589 ob- servations, atrial premature with 2,779 observations, ventricular malfunction with 7,236 observations, fusion of ventricular and normal with 803 observations and a last class of 8,039 unclassified observations. 48 half-hour excerpts of two-channel ambulatory ECG recordings. [125] [124]
Alzheimer’s Disease EEG dataset 88 participants, categorized into three groups: 36 individuals diagnosed with Alzheimer’s disease(AD group), 23 diagnosed with Frontotemporal Dementia (FTD group), and 29 healthy subjects (CN group). The duration of the disease was measured in months and the median value was 25 with IQR range (Q1-Q3) being 24 - 28.5 months. [126] [110]
Traffic Cab Trajectory datasets Real World Trajectory data of Chengdu, Harbin and Xi’an City from Didi Chuxing, around 5.6 million trajectory number. Starting from November 1, 2016, to November 30, 2016. [81] [81, 17, 65]
PeMS08 The traffic data are aggregated into every 5-minute interval from the raw data collected by the Caltrans Performance Measurement System. From July to August in 2016, containing 1979 detectors on 8 roads. [82] [59, 62, 63]
MAAD-Highway dataset A dataset for multi-agent anomaly detection based on the OpenAI Gym MultiCarRacing-v0 environment. The sequences are sub-sampled to 10 Hz with a segment length of T = 1.5 seconds. [64] [66, 67]
LargeST Traffic flow data of California, including a total number of 8,600 sensors over road networks. Starting from January 1, 2017, to December 31, 2021. [283] [283]
Recommendation Amazon-book This dataset contains product reviews and metadata from Amazon, including 142.8 million reviews. May 1996 - July 2014. $\setminus$ [132, 135]
Yelp 908,915 tips by 1,987,897 users Over 1.2 million business attributes like hours, parking, availability, and ambience Aggregated check-ins over time for each of the 131,930 businesses $\setminus$ $\setminus$ [132]
ML-IM 1 million ratings from 6000 users on 4000 movies. $\setminus$ [133] [132, 134, 95]
Amazon-beauty Over 2 Million+ customer reviews and ratings of Beauty related products sold on their website. $\setminus$ $\setminus$ [134, 36, 95]
Amazon-toys 10,000 toy products on Amazon.com $\setminus$ $\setminus$ [134, 36, 95, 135]
Climate and Weather ERA5 The fifth generation ECMWF atmospheric reanalysis of the global climate covering the period from January 1940 to present. ERA5 provides hourly estimates of a large number of atmospheric, land and oceanic climate variables. The data cover the Earth on a 31km grid and resolve the atmosphere using 137 levels from the surface up to a height of 80km. ERA5 includes information about uncertainties for all variables at reduced spatial and temporal resolutions. 1940 to present. [136] [16, 137, 84, 27]
Data from IPSL-CM5A ESM A full earth system climate. $\setminus$ [138] [15]
GOES West data Highresolution atmospheric measurements over the Pacific Ocean in near real-time with the explicit goal of improving weather forecasting capabilities. $\setminus$ $\setminus$ [27]
Energy and Electricity GEFCom2014 Load Forecasting Data $\setminus$ 7 years of matching load and temperature data [139] [107, 28]
ACN-Data Collect detailed data about each charging session in the system. $\setminus$ [140] [96]
Low Carbon London Project Data Contains electricity consumption data of 5198 customers from November 2011 to February 2014 with a time granularity of 30 minutes. November 2011 to February 2014 [141] [142]
Low Carbon London Project Data Contains electricity consumption data of 5198 customers from November 2011 to February 2014 with a time granularity of 30 minutes. November 2011 to February 2014 [141] [142]
Audio LJ Speech dataset 13,100 short audio clips of a single speaker reading passages from 7 non-fiction books. The texts were published between 1884 and 1964, and are in the public domain. The audio was recorded in 2016-17 by the LibriVox project and is also in the public domain. [68] [20, 21, 22, 91]
SC09 dataset 31,158 training utterances (8.7 hours in total) by 2,032 speakers. $\setminus$ [273] [91, 274]
@@
In detail, we will discuss the models in different applications, aiming to show light on the model design of practical applications. Besides, the main datasets in each application domains are shown in Table II , offering convenience for future research.
7.1 Healthcare
In recent years, diffusion models have emerged as a powerful class of generative models with significant implications for time series in healthcare. These models, known for their ability to generate high-fidelity samples through a process of gradual refinement, have found diverse applications ranging from the synthesis of electronic health records (EHRs) to the enhancement of biomedical signal analysis. The following are some main topics in this area:.
  • EEG Signal Synthesis and Enhancement: [74] introduces a generalized probabilistic diffusion model tailored for the synthesis of electrocardiogram (ECG) signals, aiming to support cardiac health research and diagnostic training with realistic data generation. [128] demonstrate the potential of diffusion models to create synthetic EEG signals, facilitating sleep disorder studies and neurological research. [57, 123] focus on enhancing the quality of ECG signals, addressing issues like baseline wander and noise, and generating conditional ECG signals based on specific patient states.
  • Healthcare Data Augmentation and Synthesis: [18] highlights the use of diffusion models to augment datasets for seizure forecasting, thereby improving the robustness and accuracy of predictive models. [30, 31] explore the synthesis of realistic EHRs, tackling the challenges of data privacy and scarcity by providing synthetic datasets for research and model training.
  • Diagnostic and Predictive Analytics: [50] leverages generative models to enhance the detection of autism, showcasing the broad applicability of these models in diagnosing and understanding complex conditions. In [49], diffusion models are applied to forecast critical physiological parameters in ICUs, demonstrating their potential to save lives by predicting adverse events before they occur.
  • Novel Methodologies and Techniques: [58] exemplifies the versatility of diffusion models in handling multiple tasks simultaneously, such as predicting mortality in critically ill patients, showcasing the potential for comprehensive care management systems.
7.2 Traffic
@@
There are also many works for traffic applications using diffusion models. The following are some of them. DiffUFlow [35] employs a denoising diffusion model to enable the model to capture the stochastic nature of urban flows and generate accurate and fine-grained forecasting. DiffTraj [81] presents a groundbreaking approach to generating GPS trajectories using a spatial-temporal diffusion probabilistic model. The model works by reconstructing geographic trajectories from white noise through a reverse trajectory denoising process, effectively turning random noise into meaningful trajectory data that reflects real-world movement patterns. DVGNN architecture [59] comprises an encoder component that learns latent node embeddings through GCN layers and a decoder component that treats the relationships of latent states as a diffusion process subject to stochastic differential equations. This allows for the inference of internal causal relationships among neighbor nodes and the formation of dynamic causal graphs. [61] addresses the growing risk of collisions between space objects by developing a diffusion model that predicts the positional uncertainty of objects during close encounters. [83] introduces a model that utilizes denoising diffusion probabilistic models for identifying anomalies in vehicle trajectories, aimed at improving the detection of unusual patterns in vehicle movements. SpecSTG [88] aims to efficiently handle the complexities of spatio-temporal data, offering a probabilistic approach to forecasting traffic flows. This method is distinguished by its speed and accuracy in generating forecasts and is a promising solution for real-time traffic management and planning applications.
7.3 Sequential Recommendation

### v5

@@ 6. Data Perspective
In this section, we analyze current work from a data perspective, including time series and spatio-temporal data, as shown in Figure 4 . We focus on introducing how diffusion models capture the unique properties of different data modalities.
6.1. Time Series
@@
6.1.1. Univariate Time Series
@@
Univariate time series is characterized by having only one variable of interest observed over a period of time. Data in this category includes ECG signal, audio, electricity load, etc. The complex sequential patterns (e.g., the trend and the periodicity) are the core data property in univariate time series, which pose great challenges in developing relevant diffusion models.
@@
Diffusion models for univariate time series are primarily developed to model the uncertainty present in the data, facilitating tasks such as probabilistic forecasting or data generation. An essential element within the diffusion model is the denoising network, which determines the extent of noise removal at each stage. In the context of univariate time series, the denoising network commonly employs the 1-D CNN to identify sequential patterns in the input data. For instance, both WaveGrad (Chen et al., 2020) and DiffWave (Kong et al., 2020) incorporate a 1-D CNN into the denoising network to extract local sequential features from audio sequences. Furthermore, some studies opt to transform the univariate time series into the frequency domain initially to enhance the capture of long-term sequential correlations from a global standpoint. This transformation converts the time series from 1-D to 2-D data, enabling the use of a 2-D CNN in the denoising network to capture correlations within the frequency domain, such as (Koa et al., 2023) for the stock price forecasting and DiffLoad (Wang et al., 2023b) for electricity load forecasting.
6.1.2. Multivariate Time Series
@@
Multivariate time series is characterized by having multiple variables of interest observed over the same period of time. Instead of the sequential patterns in each variate, multivariate time series also leverage the interdependencies among different variables to capture more comprehensive information for downstream tasks.
@@
Beyond approaches that focus on univariate time series, there have been efforts towards multivariate time series. Multivariate time series is naturally a 2-D data similar to image data, where the signal of each variate can be considered as one channel in the image. To this end, diffusion models for multivariate time series usually adopt the vanilla U-Net as the denoising net, which is a common practice in diffusion models for images.
@@
For example, (Durall et al., 2023) first transforms demultiplexing, denoising, and interpolation into image-to-image transformation tasks, then leverages the diffusion model for different tasks. In the demultiplex task, the diffusion model learns to remove the multiples without removing the primary energy. In the denoising and interpolation task, the diffusion model learns to eliminate undesired uncorrelated noise and recover the missing values, while preserving the inherent characteristics of the data. MIDM (Wang et al., 2023d) proposes a novel multivariate imputation diffusion model that incorporates correlations between observed and missing values. By re-deriving the ELBO of the conditional diffusion model, MIDM ensures consistency between observed and missing data, thus leading to improved imputation accuracy. Diff-E (Kim et al., 2023b) further utilizes the diffusion model for representation learning, which combines DDPMs with a conditional autoencoder to enhance the decoding performance of speech-related EEG signals.
6.2. Spatio-Temporal Data
@@
In real-world systems, a myriad of elements interact with each other both spatially and temporally, resulting in a spatio-temporal composition. Spatio-temporal data (STD) is the de facto most popular data structure for injecting such structural information into the formulation of practical problems. In this section, we introduce developments of diffusion models for spatio-temporal data, mainly with two data modalities: (1) the Spatio-Temporal Graph (STG) and (2) the Spatio-Temporal Trajectory (STT). Unlike diffusion models for time series data, existing works in STD need to model the temporal dependencies as well as the spatial dependencies, making the task even more challenging.
6.2.1. Spatio-Temporal Graph
@@
The spatio-temporal graph is generated by sensors at different places in a period of time, where the correlation of those sensors is typically described as a graph. We categorize diffusion models for spatio-temporal graphs into domain-oriented and domain-agnostic works.
@@
Most of the existing works fall into the category of domain-oriented spatio-temporal graph diffusion models. They typically leverage the powerful distribution learning abilities of diffusion models for spatio-temporal data mining tasks in specific domains, such as traffic and climate. In this research trend, the majority of works come from the traffic domain. DiffUFlow (Zheng et al., 2023) represents pioneering efforts as traffic diffusion models, which convert fine-grained urban flow inference into a denoising diffusion process. SpecSTG (Lin et al., 2024) advances the field by conducting the diffusion process in the spectral space projected by the STG, resulting in faster inference speeds. In a similar vein, (Liang et al., 2023) uses a diffusion model to infer the link probability and reconstruct causal graphs in the decoder stage adaptively for the STG forecasting task. Furthermore, (Lin et al., 2023b) solves the origin-destination travel time estimation problem with a two-step process, where the first step predicts the possible travel route with a diffusion model conditioned on the given origin-destination pair. Beyond the traffic domain, there have been initiatives towards the climate. A notable example is (Chen et al., 2024b), which quantifies the uncertainty of air quality forecasting based on the spatio-temporal diffusion model. Similarly, SRNDiff (Ling et al., 2024) conducts the short-term rainfall nowcasting through the conditional diffusion model.
@@
In the second group, researchers focus on developing domain-agnostic models, which can achieve promising performance across a variety of domains. For instance, DiffSTG represents a pioneering effort by introducing a unified diffusion framework for multiple STG tasks, such as forecasting and imputation. Concurrently, PriSTI (Liu et al., 2023c) and (Yun et al., 2023) study the spatio-temporal imputation task while treating the task of recovering unobserved values as a denoising process conditioned on the observed values. Similarly, DYffusion (Cachay et al., 2023) models the temporal dynamics directly within diffusion steps, leading to a stochastic, time-conditioned forecasting network. Another noteworthy contribution is (Hu et al., 2023) under the domain-agnostic and task-agnostic setting, which aims to harness the power of diffusion models, and unify diffusion models for probabilistic spatio-temporal graph learning.
6.2.2. Spatio-Temporal Trajectory
@@
Spatio-temporal trajectory is a sequence of locations ordered by time that describes the movements of an object in a geographical space. Analysis of spatio-temporal modality is particularly crucial to discovering the mobility patterns of moving objects, which serves as the foundation for many downstream tasks, such as POI recommendation and next-location forecasting. In terms of diffusion models for trajectory data, most works are developed to solve the trajectory generation task. They leverage the ability of the diffusion model to learn high-dimensional data distribution while injecting spatio-temporal correlation into the diffusion process. As exemplified by DiffTraj (Zhu et al., 2024a), which generates the GPS trajectory with the diffusion probabilistic model in an unconditioned manner. Similarly, (Chu et al., 2024) utilizes the diffusion model for generating high-quality human mobility data. More recently, Diff-RNTraj (Wei et al., 2024) further generates the trajectory constrained by the road-network. Beyond generation, there is a growing interest in applying diffusion models for trajectory forecasting tasks, and (Yao et al., 2023a) and (Liu et al., 2024a) exemplify this trend.
7. Application Perspective
@@
In this section, we summarize and discuss the most important applications of diffusion models in time series and spatio-temporal data, including healthcare, smart city, recommendation, climate and weather, energy and electricity, audio, video, and so on. The toy signal examples of each application are shown in  Figure 7 .
@@
Figure 7. Examples of time series and spatio-temporal data for different application scenarios.
@@
In detail, we will discuss the models in different applications, aiming to show light on the model design of practical applications.
7.1. Healthcare
In recent years, diffusion models have emerged as a powerful class of generative models with significant implications for time series in healthcare. These models, known for their ability to generate high-fidelity samples through a process of gradual refinement, have found diverse applications ranging from the synthesis of electronic health records (EHRs) to the enhancement of biomedical signal analysis. The following are some main topics in this area. EEG Signal Synthesis and Enhancement: (Neifar et al., 2023) introduces a generalized probabilistic diffusion model tailored for the synthesis of electrocardiogram (ECG) signals, aiming to support cardiac health research and diagnostic training with realistic data generation. (Aristimunha et al., 2023) demonstrates the potential of diffusion models to create synthetic EEG signals, facilitating sleep disorder studies and neurological research. (Li et al., 2023b; Alcaraz and Strodthoff, 2023) focus on enhancing the quality of ECG signals, addressing issues like baseline wander and noise, and generating conditional ECG signals based on specific patient states. Healthcare Data Augmentation and Synthesis: (Shu et al., 2023) highlights the use of diffusion models to augment datasets for seizure forecasting, thereby improving the robustness and accuracy of predictive models. (Yuan et al., 2023c; Ceritli et al., 2023) explore the synthesis of realistic EHRs, tackling the challenges of data privacy and scarcity by providing synthetic datasets for research and model training. Diagnostic and Predictive Analytics: (Li et al., 2023c) leverages generative models to enhance the detection of autism, showcasing the broad applicability of these models in diagnosing and understanding complex conditions. In (Chang et al., 2024), diffusion models are applied to forecast critical physiological parameters in ICUs, demonstrating their potential to save lives by predicting adverse events before they occur. Novel Methodologies and Techniques: (Zhao et al., 2024b) exemplifies the versatility of diffusion models in handling multiple tasks simultaneously, such as predicting mortality in critically ill patients, showcasing the potential of comprehensive care management systems.
7.2. Traffic
@@
There are also many works for traffic applications using diffusion models. The following are some of them. DiffUFlow (Zheng et al., 2023) employs a denoising diffusion model to enable the model to capture the stochastic nature of urban flows and generate accurate and fine-grained forecasting. DiffTraj (Zhu et al., 2024a) presents a groundbreaking approach to generating GPS trajectories using a spatio-temporal diffusion probabilistic model. The model works by reconstructing geographic trajectories from white noise through a reverse trajectory denoising process, effectively turning random noise into meaningful trajectory data that reflects real-world movement patterns. DVGNN architecture (Liang et al., 2023) comprises an encoder component that learns latent node embeddings through GCN layers and a decoder component that treats the relationships of latent states as a diffusion process subject to stochastic differential equations. This allows for the inference of internal causal relationships among neighbor nodes and the formation of dynamic causal graphs. (Guimarães et al., 2023) addresses the growing risk of collisions between space objects by developing a diffusion model that predicts the positional uncertainty of objects during close encounters. (Li et al., 2024a) introduces a model that utilizes denoising diffusion probabilistic models for identifying anomalies in vehicle trajectories, aimed at improving the detection of unusual patterns in vehicle movements. SpecSTG (Lin et al., 2024) aims to efficiently handle the complexities of spatio-temporal data, offering a probabilistic approach to forecasting traffic flows. This method is distinguished by its speed and accuracy in generating forecasts and is a promising solution for real-time traffic management and planning applications.
7.3. Sequential Recommendation